OpenStudy (sbuck98):

Can someone show me how to solve something using the addition method? I'm not sure how to do it, I know how to use the substitution method but they're two different things.

OpenStudy (sbuck98):

x-2y=0 5x+2y=24

OpenStudy (tootzrll):

the addition method is any number taking on more numbers (adding) and concluding with a total (or sum) so 7 + 8 = 15 and 15 is the sum

OpenStudy (sbuck98):

So how would I do the one above?

OpenStudy (tootzrll):

take x-2y=0 and flip it. 2y-0=x

OpenStudy (sbuck98):

But, how do I solve it, like I don't understand the process.

OpenStudy (tootzrll):

since 2y - 0 is 2y, answer it as 2y=x. now substitute 2y everywhere u see x. so ur equation now is 5(2y) + 2y = 24 do u hav it from here or do u need more help?

OpenStudy (sbuck98):

Yes:) Thank you so much

OpenStudy (tootzrll):

u are very welcome :)) tag me next time u need help